OpenStudy (anonymous):

Dose anyone know what (z) = In z-8/6=-4

OpenStudy (allank):

hint: if ln a = b, then e^b=a

OpenStudy (anonymous):

i don't get it?

OpenStudy (allank):

To solve this, you need to understand what 'e' is and how it relates to ln. Are you familiar with that?

OpenStudy (anonymous):

where do u get e from?

OpenStudy (allank):

The value of 'e' is pre-loaded into your calculator.

OpenStudy (allank):

Back to my question: Do you know how 'ln' and 'e' are related?

OpenStudy (anonymous):

im asking what z is ?

OpenStudy (allank):

Yes, in order to find the value of z, you need the knowledge of the relationship between ln and e. So that when I say that In z-8/6=-4 can be written as e^-4=z-8/6 You understand why. From that step, e^-4=z-8/6 , finding the value of z is just one step away.

OpenStudy (allank):

Kindly read about the relationship between 'ln' and 'e' so that you understand how I got to e^-4=z-8/6 Then solve for z from the above equation. Best wishes.

Directrix (directrix):

Is this expression [ In z-8/6=-4 ] supposed to be [ log z - 8] / 6 = -4 or perhaps [ log (z - 8) ] / 6 = -4 ?